# AI-Generated Compromises for Coalition Formation: Modeling, Simulation, and a Textual Case Study

### v1

@@ 1 Introduction
Although Elkind et al. [10] discuss the theoretical dynamics of forming coalitions via compromise proposals, they do not specify how such proposals are generated. We address this gap by introducing the notion of an AI mediator that systematically produces compromise documents. Specifically, we embed potential texts in a semantic metric space and employ modern natural language processing (NLP) techniques to measure distances between documents, thereby identifying compromise points that better align with multiple agents’ preferences. This approach is particularly relevant for large-scale, democratic text editing tasks—such as drafting a constitution for a decentralized autonomous organization (DAO) [14]—which existing collaborative platforms (e.g., Google Docs, Notion, Wikipedia) do not address in a structured, consensus-driven manner. By modeling documents as points in a high-dimensional embedding space, the AI mediator can propose new drafts that balance diverse viewpoints, thus paving the way for a more democratic editing process.
@@ 1.1 Model State
The model is defined by the following fixed components [^1] This is a centralized description for ease of presentation. Conceptually, we envision a decentralized setting, where the AI-mediator operates as a non-centralized tool available to individual coalitions. That is, coalitions may grow in a bottom-up manner, each using an instance of the mediator independently.:
@@
    We assume $z\in[n]:=\{1,\ldots,n\}$; that is, the number of coalitions does not exceed the number of agents. The notation $[z]$ refers to the index set of the current coalition structure.
@@ 1.2 Initialization, Iterative Process, and Halting Conditions
The model contains an entity—the AI-mediator—which is the workhorse of the process.
Definition 1 (AI-mediator).
An AI-mediator $M$ is a function that gets as input a coalition structure $D$ and returns as output a tuple $(d_{i},d_{j},p)$ with $d_{i},d_{j}\in D$ and $p$ a point in the metric space.
Intuitively, the AI-mediator suggests that two coalitions, $d_{i}$ and $d_{j}$, merge around a compromise point $p$. Given the current coalition structure $D$, the mediator returns a triple $(d_{i},d_{j},p)$, where $p$ is proposed as a new joint position.
@@
A constitution $const$ gets as input a tuple $(d_{i},d_{j},p)$ by the AI-mediator and, when applied on $d_{i}$ – and based on the votes of the agents in $d_{i}$, as described by $\{vote(v,p):v\in d_{i}\}$ – returns an assignment to a coalition for each $v\in d_{i}$, namely $const(v)\in\{d_{i},d^{p}\}$, where $d^{p}$ describes the coalition to be possibly-formed around the suggested compromise point $p$.
@@ 2 Concrete Model Realizations
We provide concrete realizations of the following ingredients: agent models (in Section 2.1), coalition constitutions (in Section 2.2), and AI-mediators (in Section 2.3). These concrete realizations are used later, for the 2D Euclidean setting presented in the appendix and the setting that involves text documents. Furthermore, some of the details next are needed for the computer-based simulations that follow.  [^3] We consider realizations of the model in which all agents share the same agent model; all coalitions share the same constitution; and there is one AI-mediator throughout the process. We discuss other options in Section 5.
@@ 2.1 One Concrete Agent Model
Specifically, to model different types of non-deterministic agents, we introduce a parameter $\sigma\geq 0$, where larger values of $\sigma$ results in a more altruistic agent behavior as compared to the simplest agent model described above. Mathematically, we use a half (positive) Gaussian distribution to “enlarge” a bit the region for which the agent approves the proposal (i.e., so that an agent will approve a proposal even if it is farther away from its ideal point, compared to the status quo; but with diminishing probability of doing so); formally, we have the following definition of $F$ (note that the else case is 00 in case of $\sigma=0$):
@@ 2.2 Two Concrete Constitutions
    $\text{if }|\{v\in C_{i}:\text{vote}(v,p)=1\}|\geq Q,\text{ then }\text{const}(v):=\text{vote}(v,p);$
    otherwise, $\text{const}(v):=0$. Agents who disapprove remain in $d_{i}$. [^5] Our model builds upon and generalizes aspects of the framework proposed by Elkind et al. In the case of coalition discipline with deterministic agents and unanimous approval ($Q=|C_{i}|$), we recover their convergence results under the constraint that agents only transition to strictly preferred larger coalitions.
  • No Coalition Discipline is a special case of the above with $Q=0$, where each agent independently decides whether to join the new coalition: $\text{const}(v):=\text{vote}(v,p)$.
@@
2.3 Several Concrete AI-Mediators
Recall that an AI-mediator takes as input a coalition structure $D$ and returns two coalitions, $d_{i}$ and $d_{j}$, and a compromise point $p$. It is convenient to break the description of our realizations into the two main tasks of AI-mediators, namely: (1) choosing the coalitions $d_{i}$ and $d_{j}$ to suggest $p$ to; and (2) choosing the compromise point $p$ to suggest to $d_{i}$ and $d_{j}$.
@@ 2.3 Several Concrete AI-Mediators
Our AI-mediators proceed by first computing the centroid of the coalitions’ ideal points, weighted by the coalition sizes. Formally: $centroi\mathbf{d}(D)=\arg\min_{x\in X}\frac{1}{n}\cdot\sum_{i\in|D|}|C_{i}|\cdot\mathbf{d}(x,p_{i})$. Using the centroid, the AI-mediators consider the distance of each coalition from the centroid, denoted by $\mathbf{d}(p_{i},centroid)$. The selection process is guided by a parameter $\alpha\in[-1,1]$, intuitively ranging from whether the closest coalitions to the centroid are preferred ($\alpha=-1$), the furthest ones are preferred ($\alpha=1$), or there is no significance ($\alpha=0$) to the distance from the centroid.
@@
Subsequently, the AI-mediator assigns a probability $prob(d_{i})$ to each coalition $d_{i}$, proportionate to its scores: $prob(d_{i})=\frac{S_{i}}{\sum_{i=1}^{|D|}S_{i}}$. In practice, the AI-mediator probabilistically chooses one coalition based on these probabilities and then selects the closest coalition to the initially chosen one.
@@
Given coalitions $d_{i}=(C_{i},p_{i})$ and $d_{j}=(C_{j},p_{j})$, the AI-mediator selects a compromise point $p\in X$ that minimizes the weighted sum of distances to $p_{i}$ and $p_{j}$, with weights proportional to coalition sizes:
@@
The AI-mediator is assumed to know all agents’ ideal points. This may result from voluntary disclosure or be treated as a modeling assumption. While this enables computation of globally optimal coalitions under a defined objective, the current mediator applies local, myopic merges. Designing optimal, forward-looking mediators is left for future work.
@@ 3 Related Work
Coalition formation in a metric space has been studied from a multiagent system context [6, 37, 31]. We build upon the theoretical framework of Elkind et al. [10], which introduced a model for deliberative coalition formation in metric spaces. Their work presents a transition system to capture the dynamics of the coalition formation process. While Elkind et al. describe the formation process in detail, they assume that compromise points are provided by an external source (an oracle), without specifying how these points should be determined. Our contribution addresses this gap by introducing AI-mediators that algorithmically suggest compromise points, allowing coalitions to unite around majority-supported proposals. We implement and optimize these AI-mediators to make the coalition formation process both practical and efficient. We also demonstrate that, under a specific configuration of our model, it aligns with Elkind’s model, thereby inheriting their theoretical results for that configuration thus our model generalizes Elkind’s model. For the general case we show simulations that show convergence rates are very good even for large instances. In developing AI-mediators, we utilize NLP techniques and LLMs; this relates to NLP-based recommendation systems [3, 33], where models suggest content based on user preferences, and to recent work in Generative Social Choice [13], which explores the use of LLMs to generate representative statements for social choice tasks. However, our work differs by focusing on identifying compromise points in the coalition formation process, where the goal is to find a majority-supported text or proposal. Another relevant line of work is by Bakker et al. [2], who study how machines can assist in finding agreements among individuals with diverse preferences. Their approach fine-tunes LLMs to generate statements that maximize the expected approval of a group, which is conceptually similar to our use of AI for proposing compromise points. However, our model incorporates an iterative coalition formation process, making it distinct in its operational dynamics. We also mention Yang et al. [35] that investigate how GPT-4 and LLaMA-2 behave in voting scenarios compared to human voters. They show that voting methods, presentation order, and temperature settings can significantly influence LLM choices, often reducing preference diversity and risking bias.
@@
In the context of Dynamic Coalition Formation, there is significant prior work on how agents with diverse preferences form and adapt coalitions to achieve consensus [32, 20]. This is relevant, as coalition formation plays a key role in decision-making processes, especially when agents aim to form majority-supported agreements [27]. Our approach to coalition formation in metric spaces also draws on existing research in spatial voting models [11, 16]. We are also motivated by psychological research on the ability of agents to objectively evaluate proposals. Mikhaylovskaya et al. [24] provide evidence that AI-based mediators can mitigate human biases, making AI a promising tool for generating neutral, data-driven compromise points. This motivates our use of AI-mediated coalition formation, where agents can evaluate AI-suggested compromise points to find collective agreements. Our work also draws inspiration from negotiation-based approaches to coalition formation [26, 15, 29, 10, 17, 34, 4, 12], which offer valuable insights into how agents with divergent preferences negotiate and form coalitions. These approaches further reinforce the relevance of AI mediation in improving the efficiency and effectiveness of coalition formation in multi-agent systems.
4 AI-Mediators in a Textual Space
@@
As our ultimate goal relates to text aggregation – i.e., to enable an agent community to converge towards a majority-supported textual document. So, we wish to utilize the AI-mediators framework (demonstrated also in a Euclidean space in the appendix) to a setting in which the metric space contains textual documents and coalitions form around different texts, until a majority-supported textual document is identified. We describe our specific model; and then report on computer-based simulations.
4.1 Modeling AI-Mediators in a Textual Space
@@
Our general solution works as follows. We use word embedding (a standard NLP technique) to translate texts into numerical-valued vectors; this is crucial as, after applying such a word embedding, we are then able to compute distances between the embedded coalition ideal points and use the AI-mediators of the Euclidean space. In this work, we use Google’s Universal Sentence Encoder [7]: this is a pre-trained model that converts sentences into fixed-size vectors, capturing their semantic meanings. (The Universal Sentence Encoder is designed to generate 512-dimensional embedding vectors, providing a semantic representation of sentences.) Thus our embedded metric space contains as elements all those $512$-length vectors that can be the output of the Universal Sentence Encoder. As a distance measure in this space, we use a commonly-used metric in NLP: the squared cosine distance [30]. [^6] It is indeed a metric: $d_{\text{sqrt cosine}}(A,B):=\sqrt{2-2\cdot\text{similarity}(A,B)}:=\sqrt{2-2\cdot\frac{A\cdot B}{\|A\|\cdot\|B\|}}\in[0,2]$, where $A$ and $B$ are two (embedded) vectors, and $\text{similarity}(A,B)$ measures semantic similarity of the vectors: $\frac{A\cdot B}{\|A\|\cdot\|B\|}\in[-1,1].$ The AI mediator guides the coalition formation by proposing sentences to two coalitions within the given word limit (in our simulations, $15$ words). In each iteration, the mediator’s goal is to find two coalitions to suggest a sentence that minimizes the squared cosine distance between the embedding vector of the chosen sentence and the weighted average of the embedding vectors of the two coalition points.
@@
Our approach to coalition formation in the domain of text relies on the integration of OpenAI’s GPT-3.5-turbo-1106 model (https://openai.com/blog/chatgpt) with a temperature parameter (responsible for the randomness of results) of $0.75$ as recommended in some of the documentation to provide a good trade-off for applications like ours where the output should be coherent but still allow for some diversity and creativity. This LLM takes $3$ key roles within our framework:
@@ 4.2 Simulations-Based Analysis
  • Coalition Formation Process: We employed an iterative pursuit in an embedding space using squared cosine distance; the prompt given to GPT was: “Give me T different sentences that are well structured about how to deal with Y with at most of 15 words” (T being the number of agents, and Y being any topic – global warming in our case); to initialize the singleton coalition with introducing noise ($I=True$), the LLM was requested to provide a sentence resembling the ideal sentence of each agent, rather than introducing additional noise through a normal distribution as conducted in the euclidean case presented in the appendix. These function as the singleton coalition sentences to be embedded into the Euclidean space. The prompt given to the GPT was: “Give me a well-structured sentence with a maximum of 15 words, resembling this sentence: Z” (where Z represents an ideal sentence of an agent).
  • Sentence Selection Process: For each proposed sentence to the two coalitions, the LLM was tasked with generating $10$ sentences that effectively combined both coalition sentences. We followed best practices for structured prompt design and multi-step reasoning [19], including these concepts:
@@
    We used the following prompts and messages given to GPT 3.5 ($5$ options in total):
    • Mediator 1: Prompt: Generate $10$ possible different well-structured sentences that aggregate the following two sentences. Make sure each sentence has at most 15 words. Number your answers (i.e., 1), 2), 3), 4), 5), and so on) for each sentence you propose. Message: You are a mediator trying to find agreed wording for how to deal with global warming based on existing sentences. Give a straightforward answer with no introduction to help people reach an agreed wording of a coherent sentence. (The proposed sentence was selected based on the minimal squared cosine distance between its embedding and the weighted average embedding vector, considering the two embedding vectors of the coalitions and their sizes.)
@@
Table 1: A comparison of different AI mediators-each corresponding to different prompts and LLM-usage strategy. The mean number of iterations until convergence is shown, validated with $95\%$ confidence using ANOVA and Tukey HSD.

### v2

@@ 1 Introduction
Although Elkind et al. [10] discuss the theoretical dynamics of forming coalitions via compromise proposals, they do not specify how such proposals are generated. We address this gap by introducing the notion of an mediator that systematically produces compromise documents. Specifically, we embed potential texts in a semantic metric space and employ modern natural language processing (NLP) techniques to measure distances between documents, thereby identifying compromise points that better align with multiple agents’ preferences. This approach is particularly relevant for large-scale, democratic text editing tasks—such as drafting a constitution for a decentralized autonomous organization (DAO) [14]—which existing collaborative platforms (e.g., Google Docs, Notion, Wikipedia) do not address in a structured, consensus-driven manner. By modeling documents as points in a high-dimensional embedding space, the mediator can propose new drafts that balance diverse viewpoints, thus paving the way for a more democratic editing process.
@@ 1.1 Model State
The model is defined by the following fixed components [^1] This is a centralized description for ease of presentation. Conceptually, we envision a decentralized setting, where the mediator operates as a non-centralized tool available to individual coalitions. That is, coalitions may grow in a bottom-up manner, each using an instance of the mediator independently.:
@@
[…]
@@ 1.2 Initialization, Iterative Process, and Halting Conditions
The model contains an entity—the mediator—which is the workhorse of the process.
Definition 1 (mediator).
A mediator $M$ is a function that gets as input a coalition structure $D$ and returns as output a tuple $(d_{i},d_{j},p)$ with $d_{i},d_{j}\in D$ and $p$ a point in the metric space.
Intuitively, the mediator suggests that two coalitions, $d_{i}$ and $d_{j}$, merge around a compromise point $p$. Given the current coalition structure $D$, the mediator returns a triple $(d_{i},d_{j},p)$, where $p$ is proposed as a new joint position.
@@
A constitution $const$ gets as input a tuple $(d_{i},d_{j},p)$ by the mediator and, when applied on $d_{i}$ – and based on the votes of the agents in $d_{i}$, as described by $\{vote(v,p):v\in d_{i}\}$ – returns an assignment to a coalition for each $v\in d_{i}$, namely $const(v)\in\{d_{i},d^{p}\}$, where $d^{p}$ describes the coalition to be possibly-formed around the suggested compromise point $p$.
@@ 2 Concrete Model Realizations
We provide concrete realizations of the following ingredients: agent models (in Section 2.1), coalition constitutions (in Section 2.2), and mediators (in Section 2.3). These concrete realizations are used later, for the 2D Euclidean setting presented in the appendix and the setting that involves text documents. Furthermore, some of the details next are needed for the computer-based simulations that follow.  [^3] We consider realizations of the model in which all agents share the same agent model; all coalitions share the same constitution; and there is one mediator throughout the process. We discuss other options in Section 5.
@@ 2.1 One Concrete Agent Model
Specifically, to model different types of non-deterministic agents, we introduce a parameter $\sigma\geq 0$, where larger values of $\sigma$ results in a more altruistic agent behavior as compared to the simplest agent model described above. Mathematically, we use a half (positive) Gaussian distribution to “enlarge” a bit the region for which the agent approves the proposal (i.e., so that an agent will approve a proposal even if it is farther away from its ideal point, compared to the status quo; but with diminishing probability of doing so); formally, we have the following definition of $F$ (note that the else case is 0 in case of $\sigma=0$):
@@ 2.2 Two Concrete Constitutions
    $\text{if }|\{v\in C_{i}:\text{vote}(v,p)=1\}|\geq Q,\text{ then for each }v\in C_{i}:\quad\text{const}(v):=\begin{cases}d_{p}&\text{if }\text{vote}(v,p)=1\\ d_{i}&\text{otherwise}\end{cases}$
    otherwise, $\text{const}(v):=d_{i}$ for all $v\in C_{i}$. [^5] Our model builds upon and generalizes aspects of the framework proposed by Elkind et al. In the case of coalition discipline with deterministic agents and unanimous approval ($Q=|C_{i}|$), we recover their convergence results under the constraint that agents only transition to strictly preferred larger coalitions.
  • No Coalition Discipline is a special case of the above with $Q=0$, where each agent independently decides whether to join the new coalition:
    $\text{const}(v):=\begin{cases}d_{p}&\text{if }\text{vote}(v,p)=1\\ d_{i}&\text{otherwise}\end{cases}\quad\text{for each }v\in C_{i}.$
@@
2.3 Several Concrete Mediators
Recall that an mediator takes as input a coalition structure $D$ and returns two coalitions, $d_{i}$ and $d_{j}$, and a compromise point $p$. It is convenient to break the description of our realizations into the two main tasks of mediators, namely: (1) choosing the coalitions $d_{i}$ and $d_{j}$ to suggest $p$ to; and (2) choosing the compromise point $p$ to suggest to $d_{i}$ and $d_{j}$. Note that the role of the AI in our design is in the implementation of such mediators.
@@ 2.3 Several Concrete Mediators
Our mediators proceed by first computing the centroid of the coalitions’ ideal points, weighted by the coalition sizes. Formally: $centroi\mathbf{d}(D)=\arg\min_{x\in X}\frac{1}{n}\cdot\sum_{i\in|D|}|C_{i}|\cdot\mathbf{d}(x,p_{i})$. Using the centroid, the mediators consider the distance of each coalition from the centroid, denoted by $\mathbf{d}(p_{i},centroid)$. The selection process is guided by a parameter $\alpha\in[-1,1]$, intuitively ranging from whether the closest coalitions to the centroid are preferred ($\alpha=-1$), the furthest ones are preferred ($\alpha=1$), or there is no significance ($\alpha=0$) to the distance from the centroid.
@@
Subsequently, the mediator assigns a probability $prob(d_{i})$ to each coalition $d_{i}$, proportionate to its scores: $prob(d_{i})=\frac{S_{i}}{\sum_{i=1}^{|D|}S_{i}}$. In practice, the mediator probabilistically chooses one coalition based on these probabilities and then selects the closest coalition to the initially chosen one.
@@
Given coalitions $d_{i}=(C_{i},p_{i})$ and $d_{j}=(C_{j},p_{j})$, the mediator selects a compromise point $p\in X$ that minimizes the weighted sum of distances to $p_{i}$ and $p_{j}$, with weights proportional to coalition sizes:
@@
The mediator is assumed to know all agents’ ideal points. This may result from voluntary disclosure or be treated as a modeling assumption. While this enables computation of globally optimal coalitions under a defined objective, the current mediator applies local, myopic merges. Designing optimal, forward-looking mediators is left for future work.
@@ 3 Related Work
Coalition formation in a metric space has been studied from a multiagent system context [6, 37, 31]. We build upon the theoretical framework of Elkind et al. [10], which introduced a model for deliberative coalition formation in metric spaces. Their work presents a transition system to capture the dynamics of the coalition formation process. While Elkind et al. describe the formation process in detail, they assume that compromise points are provided by an external source (an oracle), without specifying how these points should be determined. Our contribution addresses this gap by introducing mediators that algorithmically suggest compromise points, allowing coalitions to unite around majority-supported proposals. We implement and optimize these mediators to make the coalition formation process both practical and efficient. We also demonstrate that, under a specific configuration of our model, it aligns with Elkind’s model, thereby inheriting their theoretical results for that configuration thus our model generalizes Elkind’s model. For the general case we show simulations that show convergence rates are very good even for large instances. In developing mediators, we utilize NLP techniques and LLMs; this relates to NLP-based recommendation systems [3, 33], where models suggest content based on user preferences, and to recent work in Generative Social Choice [13], which explores the use of LLMs to generate representative statements for social choice tasks. However, our work differs by focusing on identifying compromise points in the coalition formation process, where the goal is to find a majority-supported text or proposal. Another relevant line of work is by Bakker et al. [2], who study how machines can assist in finding agreements among individuals with diverse preferences. Their approach fine-tunes LLMs to generate statements that maximize the expected approval of a group, which is conceptually similar to our use of AI for proposing compromise points. However, our model incorporates an iterative coalition formation process, making it distinct in its operational dynamics. We also mention Yang et al. [35] that investigate how GPT-4 and LLaMA-2 behave in voting scenarios compared to human voters. They show that voting methods, presentation order, and temperature settings can significantly influence LLM choices, often reducing preference diversity and risking bias.
@@
In the context of Dynamic Coalition Formation, there is significant prior work on how agents with diverse preferences form and adapt coalitions to achieve consensus [32, 20]. This is relevant, as coalition formation plays a key role in decision-making processes, especially when agents aim to form majority-supported agreements [27]. Our approach to coalition formation in metric spaces also draws on existing research in spatial voting models [11, 16]. We are also motivated by psychological research on the ability of agents to objectively evaluate proposals. Mikhaylovskaya et al. [24] provide evidence that AI-based mediators can mitigate human biases, making AI a promising tool for generating neutral, data-driven compromise points. This motivates our use of AI-mediated coalition formation, where agents can evaluate AI-suggested compromise points to find collective agreements. Our work also draws inspiration from negotiation-based approaches to coalition formation [26, 15, 29, 10, 17, 34, 4, 12], which offer valuable insights into how agents with divergent preferences negotiate and form coalitions. These approaches further reinforce the relevance of AI based mediation in improving the efficiency and effectiveness of coalition formation in multi-agent systems.
4 Mediators in a Textual Space
@@
As our ultimate goal relates to text aggregation – i.e., to enable an agent community to converge towards a majority-supported textual document. So, we wish to utilize the mediators framework (demonstrated also in a Euclidean space in the appendix) to a setting in which the metric space contains textual documents and coalitions form around different texts, until a majority-supported textual document is identified. We describe our specific model; and then report on computer-based simulations.
4.1 Modeling Mediators in a Textual Space
@@
Our general solution works as follows. We use word embedding (a standard NLP technique) to translate texts into numerical-valued vectors; this is crucial as, after applying such a word embedding, we are then able to compute distances between the embedded coalition ideal points and use the mediators of the Euclidean space. In this work, we use Google’s Universal Sentence Encoder [7]: this is a pre-trained model that converts sentences into fixed-size vectors, capturing their semantic meanings. (The Universal Sentence Encoder is designed to generate 512-dimensional embedding vectors, providing a semantic representation of sentences.) Thus our embedded metric space contains as elements all those 512512512-length vectors that can be the output of the Universal Sentence Encoder. As a distance measure in this space, we use the square root of the cosine-based dissimilarity [30], a commonly used pseudo metric in NLP. [^6] Formally, we define $d_{\text{cos}}(A,B):=\sqrt{2-2\cdot\frac{A\cdot B}{\|A\|\cdot\|B\|}}\in[0,2],$ which corresponds to $\sqrt{2}\cdot\sin(\theta/2)$, where $\theta$ is the angle between vectors $A$ and $B$. This is a metric when restricted to the unit sphere (or more generally, to rays through the origin), but only a pseudo metric over the entire embedding space: co-linear vectors have distance zero, even if they differ in magnitude. Since our mediator constructs compromise vectors via (non-normalized) weighted averages, we do not assume that vectors lie on a sphere. The mediator guides the coalition formation by proposing sentences to two coalitions within the given word limit (in our simulations, 151515 words). In each iteration, the mediator’s goal is to find two coalitions to suggest a sentence that minimizes the squared cosine similarity pseudo metric between the embedding vector of the chosen sentence and the weighted average of the embedding vectors of the two coalition points.
@@
Our approach to coalition formation in the domain of text relies on the integration of OpenAI’s GPT-3.5-turbo-1106 model (https://openai.com/blog/chatgpt) with a temperature parameter (responsible for the randomness of results) of 0.750.750.75 as recommended in some of the documentation to provide a good trade-off for applications like ours where the output should be coherent but still allow for some diversity and creativity. This LLM takes 333 key roles within our framework:
@@ 4.2 Simulations-Based Analysis
  • Coalition Formation Process: We employed an iterative pursuit in an embedding space using squared cosine similarity pseudo metric; the prompt given to GPT was: “Give me T different sentences that are well structured about how to deal with Y with at most of 15 words” (T being the number of agents, and Y being any topic – global warming in our case); to initialize the singleton coalition with introducing noise ($I=True$), the LLM was requested to provide a sentence resembling the ideal sentence of each agent, rather than introducing additional noise through a normal distribution as conducted in the euclidean case presented in the appendix. These function as the singleton coalition sentences to be embedded into the Euclidean space. The prompt given to the GPT was: “Give me a well-structured sentence with a maximum of 15 words, resembling this sentence: Z” (where Z represents an ideal sentence of an agent).
  • Sentence Selection Process: For each proposed sentence to the two coalitions, the LLM was tasked with generating 101010 sentences that effectively combined both coalition sentences. We followed best practices for structured prompt design and multi-step reasoning [19], including these concepts:
@@
    We used the following prompts and messages given to GPT 3.5 (555 options in total):
    • Mediator 1: Prompt: Generate 101010 possible different well-structured sentences that aggregate the following two sentences. Make sure each sentence has at most 15 words. Number your answers (i.e., 1), 2), 3), 4), 5), and so on) for each sentence you propose. Message: You are a mediator trying to find agreed wording for how to deal with global warming based on existing sentences. Give a straightforward answer with no introduction to help people reach an agreed wording of a coherent sentence. (The proposed sentence was selected based on the minimal squared cosine similarity pseudo metric between its embedding and the weighted average embedding vector, considering the two embedding vectors of the coalitions and their sizes.)
@@
[…]